The angle of elevation of a tree at a place 80 m away is 12 degree 5 min.What is the height of a tree?

Grade:11

1 Answers

Vikas TU
6 years ago
theta = 12 degrees 5 mins
Height is h .
Distance is 80m
According to question a right angled triangle is formed
Tan theta  = h/d
Tan theta = h/80
0.2144 * 80 = h
h=17.152 m.
This would be the height of the tree.